The length of a rectangle is
4 meters less than twice the width. If the area of the rectangle is
286 square​ meters, find the dimensions

Answer by ikleyn(52785)

.
Let W be the width.


Then the length is L = 2W - 4,  and the area is W*L = W*(2W-4).


Your equation is


W*(2W-4) = 286,    or


W*(W-2) = 143.


You can solve this quadratic equation using the quadratic formula,

or you can mentally find decomposition of the number 143 into the product of two positive integers that differ by 2:  143 = 11*13.



Answer.  W = 13 (the width);  L = 2*W - 4 = 2*13-4 = 22 (the length).


Check.   11*18 = 286 square meters.   ! Correct !